A Food Marketing Institute found that 34% of households spend more than $125 a week on groceries. Assume the population proportion is 0.34 and a simple random sample of 147 households is selected from the population. What is the probability that the sample proportion of households spending more than $125 a week is between 0.27 and 0.43?

Note: You should carefully round any z-values you calculate to 4 decimal places to match wamap's approach and calculations. Do not use tables unless specifically told to.

Answer =  (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Using central limit theorem,

P(\hat{p} < p ) = P(Z < (\hat{p} - p) / sqrt [ p ( 1 - p) / n ]

So,

P(0.27 < \hat{p} < 0.43) = P( \hat{p} < 0.43) - P( \hat{p} < 0.27)

= P(Z < (0.43 - 0.34) / sqrt [ 0.34 ( 1 - 0.34) / 147 ) - P(Z < (0.27 - 0.34) / sqrt [ 0.34 ( 1 - 0.34) / 147 )

= P(Z < 2.3035) - P(Z < -1.7916)

= 0.9894 - 0.0366 [ From Z table ]

= 0.9528